OpenStudy (anonymous):

A pilot flies a plane south and then 600 miles west, where she lands the plane. How far south did the pilot fly the plane if she lands 610 miles from her starting point?

OpenStudy (anonymous):

just a triangle hypotenuse=610 one side =600 what is the other ? 610^2-600^2=x^2 x=110